# Debiasing Watermarks for Large Language Models via Maximal Coupling

### v1

@@ 5 Language Model Experiments
Setup.
We evaluate our watermarking method against three schemes discussed above: (1) the Gumbel-max watermark based on exponential minimal sampling by Aaronson and Kirchner (2023), (2) the soft green/red list watermark by Kirchenbauer et al. (2023a) with a moderate text distortion hyperparameter $\delta=1$, and (3) the Distribution-Preserving watermark (DiPmark) by Wu et al. (2023) with their recommended watermark hyperparameter $\alpha=0.45$. For all methods where applicable, we set $\gamma=0.5$ to control the green list size such that $|{\mathcal{G}}|=\gamma|{\mathcal{W}}|$. For a fair comparison, our scheme uses both the green tokens and the red tokens for detection as described in Section 3.2. Following Fernandez et al. (2023), for all methods, we score tokens based on the uniqueness of the full $(k+1)$-tuple and compute $p$-values using the exact null distribution rather than $z$-scores, and flag a text as watermarked if the $p$-value is lower than $\alpha=0.01$, which ensures a false positive rate to be less than $0.01$. To ensure fair comparison of text quality, we implement repeated context masking for all methods (Wu et al., 2023).
@@
To evaluate the text distortion induced by the watermark, we use the S-BERT similarity score to compare the semantic similarity between the generated text of the same model with and without watermarking (Reimers and Gurevych, 2019); the higher the similarity score, the less the distortion. We also report the true positive rate (TPR) (i.e., the fraction of correctly identifying a watermarked response among all watermarked samples) before and after the substitution attack (TPR aug.) for each scheme. Following Fernandez et al. (2023), we simulate the substitution attack by randomly and independently replacing each token with probability $0.1$. In addition, we investigate the speculative decoding setup to assess the robustness of different watermarks where the goal is to modify watermarked texts so as to improve their quality. As discussed earlier, this setup better approximates the human-language model interaction. Here, we choose Phi-3-medium-4k-instruct (14B parameters) as the target model. [^9] The source code is available at https://github.com/Xieyangxinyu/Unbiased-Watermark-via-Maximal-Coupling. We conduct comprehensive ablation studies with identical experimental configurations, also including Meta’s LLaMA model family (Dubey et al., 2024). Since both model families produce similar results, we focus on the Phi-3 family in the following sections for brevity and present the complete results in Appendix A.5.3.
@@
Table 1: Comparison of the S-BERT similarity scores and the true positive rates (TPR) among three watermarking schemes: (1) the Gumbel-max watermark (Aaronson and Kirchner, 2023), (2) the green/red list watermark (Kirchenbauer et al., 2023a), (3) the DiPmark method Wu et al. (2023) and (4) our proposed one.
Comparison against Baselines.
Table 1 reports the S-BERT similarity scores and the TPR among the four watermarking schemes. Although the Gumbel-max watermark is theoretically unbiased with the assumption of perfect pseudorandomness (Aaronson and Kirchner, 2023), experimental results show that the deterministic decoder can distort the generated text more severely than the stochastic ones. Indeed, when using the preceding $k$ tokens as the seed for the pseudorandom variables $\xi$ during decoding, perfect pseudorandomness is compromised in practice, as evidenced by the consistently lower S-BERT scores. Meanwhile, when comparing against the green/red list soft watermark and DiPmark, our method often leads to a more competitive TPR when similar S-BERT scores are achieved.
@@
Text Repetition
We quantitatively evaluate the repetitiveness issue discussed in Section 5 by analyzing the proportion of repeated $k$-token sequences in the generated text. The “Repeated” column in Table 2 reports the average percentage of times that repeated context masking is applied within each generated text, which directly reflects how frequently $k$-grams are repeated during generation. While repeated context masking was introduced by Hu et al. (2023); Wu et al. (2023); Dathathri et al. (2024) to mitigate quality degradation, our results show that the Gumbel-max approach still produces notably more repetitive text compared to other methods. For example, with $k=2$ on the FinQA dataset, Gumbel-max exhibits an 18.61% token repetition rate versus 16.77-17.80% for other methods. This suggests that in practice, deterministic dependence on prior $k$-tokens in methods like Gumbel-max can bias the generation toward repetitive patterns that persist even with repeated context masking. In contrast, our stochastic decoder’s additional randomness in sampling helps reduce such repetitive patterns while maintaining watermark effectiveness.
Speculative Decoding.
Table 3 presents a comparison of the true positive rates (TPR) between the Gumbel-max watermark (Aaronson and Kirchner, 2023) and our method under the speculative decoding setup. This setup, where a draft model proposes tokens that may be rejected by the target model, naturally mirrors how users iteratively refine generated text - making targeted, non-independent edits to improve quality rather than random token substitutions. The average rejection rates indicate the proportion of watermarked tokens that are discarded during this process. While both approaches achieve comparable TPR, the lower rejection rate of our method can translate to improved user experience - users need to make fewer revisions to achieve their desired output compared to the Gumbel-max approach.
@@
Table 3: Comparison of the true positive rates (TPR) and rejection rates between the Gumbel-max watermark (Aaronson and Kirchner, 2023) and our proposed scheme under the speculative decoding setup. By embedding a watermark only on the smaller draft model, we explore the performance of watermarking methods when subjected to targeted modifications.
Comparison of Test Statistics.
In Table 4, we experiment with the two different choices of test statistics: the sum test and HC. Our experiments confirm the insights from Section 4: the sum test consistently achieves higher true positive rates (TPR) than HC across all configurations. This aligns with our simulation findings that the sum test outperforms HC when number of tokens used to detect the watermark is relatively small ($\leq 10^{3}$). For longer generated texts, we expect HC to become increasingly competitive with the sum test, as suggested by our simulation results.
@@
6 Extension to the Soft green/red List Watermark
Lemma 6.1.
@@ 7 Conclusions
Acknowledge
This work was supported in part by NSF DMS-2310679, a Meta Faculty Research Award, and Wharton AI for Business. This work utilized resources provided by the Argonne Leadership Computing Facility at Argonne National Laboratory. We would like to thank two anonymous referees for their constructive comments that helped improve the presentation of the paper.
@@ A.1 Unbiased Watermarking Schemes and Related Works
Several alternative methods have been proposed to design provably unbiased watermarking schemes. The main idea involves modifying the decoding strategy of the language model decoder to create watermark signals without altering the token distributions. The first and widely recognized method adapts the exponential minimal sampling, as introduced by Aaronson and Kirchner (2023). This method selects the next token $w_{t}$ directly by computing $w_{t}={\mathcal{S}}(\bm{P}_{t},\zeta_{t}):=\operatorname{argmax}_{w\in{\mathcal{W}}}\log U_{w}/P_{w}$ where $\zeta_{t}=(U_{1},\ldots,U_{|{\mathcal{W}}|})$ is the random variable that consists of $|{\mathcal{W}}|$ i.i.d. copies of $U[0,1]$ and is part of the watermark key. This method is unbiased when true randomness is used for generating $\zeta_{t}$ and the proof involves the Gumbel trick (Gumbel, 1948; Fernandez et al., 2023; Li et al., 2024). However, a common practice to share the watermark key is to seed a pseudorandom number generator for $U$ with the prior $k$ tokens $w_{(t-k)},...,w_{(t-1)}$ (discussed in Section 3.3). Notice that upon fixing the $U$, the decoder is deterministic, unlike the standard decoding methods we use in this paper. When the same $k$-gram appears multiple times in the generated text, this deterministic decoder will generate the same token following these $k$-grams, leading to repetitive texts and degraded text quality (Christ et al., 2023; Kuditipudi et al., 2023). To mitigate this, Wu et al. (2023) propose to apple repeated context masking, which prevents the watermark from being applied on step $t$ if the same $k$-gram has been used to watermark previously. This method has later been adapted by Hu et al. (2023); Dathathri et al. (2024). However, the deterministic nature of this decoder can bias the generation toward repetitive patterns that persist even with repeated context masking, as we demonstrate in Section 5. Several theoretical analyses and variants of this method have been proposed (Fernandez et al., 2023; Zhao et al., 2024), but no existing work completely addresses this issue.
Another method is the inverse transform sampling strategy (Christ et al., 2023; Kuditipudi et al., 2023; Hu et al., 2023; Li et al., 2024). The idea is to first sample a standard uniform random variable $\zeta_{1:n}=(\zeta_{1},...,\zeta_{n})$, and then map $\zeta_{t}$ to the next token. Although the randomness from $\zeta_{t}$ can guarantee the unbiasedness of the token generation, designing an effective test to detect watermarks with this method is challenging: Kuditipudi et al. (2023) showed that their correlation-based detection method yields weaker power than the exponential minimal sampling method, even when perfect randomness is applied, while the detection method proposed by Hu et al. (2023) involves solving an optimization problem and requires hyperparameter tuning to achieve competitive power.
@@ 2. $H_{1}$:
$\Lambda(\zeta_{1:n})=\frac{\prod_{t=1}^{m}\mathds{1}_{\zeta_{j}\in\left[0,P_{t,{\mathcal{G}}_{t}}\right]}/P_{t,{\mathcal{G}}_{t}}}{\prod_{t=1}^{m}\mathds{1}_{\zeta_{t}\in\left[0,1\right]}}=(1-\delta)^{-m}\mathds{1}_{\zeta_{(m)}\leq 1-\delta}$
where $\mathds{1}$ denotes the indicator function. This likelihood-ratio test provides the decision rule as follows: for some critical value $c$,
@@ 1. $H_{1}^{(\mathrm{mix})}$:
Suboptimality of the Order Statistics.
@@
As observed earlier, we can assume without loss of generality that $P_{t,{\mathcal{G}}_{t}}=P_{{\mathcal{G}}}=m^{-r}$ for all $t$. By Lemma A.1, it suffices to show that the Hellinger affinity between $U[0,1]$ and $(1-\varepsilon_{m})U[0,1]+\varepsilon_{m}U[0,P_{{\mathcal{G}}}]$ behaves asymptotically as $1+o(1/m)$. Let $f(y)$ and $g(y)$ be the densities of $U[0,1]$ and $U[0,P_{{\mathcal{G}}}]$ respectively. As $f(y)=1$ and $g(y)=P_{{\mathcal{G}}}^{-1}\mathds{1}_{x\in[0,P_{{\mathcal{G}}}]}$ on $[0,1]$, the Hellinger affinity is given by
@@
$=\varepsilon_{m}^{2}\left(\mathbb{E}_{0}p(G)^{-2}\mathds{1}_{x\in[0,P_{{\mathcal{G}}}]}-1\right)$
@@
$\mathbb{E}_{0}L_{m}\to-\infty\quad\text{and}\quad\frac{\operatorname{Var}_{0}(L_{m})}{[\mathbb{E}_{0}L_{m}]^{2}}\to 0.$
@@
$\operatorname{Var}_{1}s=m\operatorname{Var}_{1}\zeta_{1}=O(m).$
@@
$F_{m}(x)=\frac{1}{m}\sum_{i=1}^{m}\mathds{1}_{\zeta_{i}\leq x},$
@@
$=\sqrt{m}\cdot\left(\frac{\frac{1}{m}\sum_{i=1}^{m}\mathbb{E}_{1}\mathds{1}_{\zeta_{i}\leq x}-x}{\sqrt{x(1-x)}}\right)$
@@
A.4 Speculative Decoding
@@
Connections to Maximal Coupling
@@
$\mathbb{P}[\texttt{Accept}]=\mathbb{P}_{w\sim\bm{Q}}[\zeta\cdot Q_{w}>P_{w}]=\sum_{w\in{\mathcal{W}}}Q_{w}\cdot\min(\frac{P_{w}}{Q_{w}},1)=\sum_{w\in{\mathcal{W}}}\min(P_{w},Q_{w})=p.$
@@
Speculative Decoding as Post-Processing
@@
$Q_{w}=\mathds{1}_{\{w=\operatorname{argmax}_{w^{\prime}\in{\mathcal{W}}}\log U_{w^{\prime}}/\tilde{P}_{w}^{\prime}\}}$
@@
$Q_{w}=\mathds{1}_{\{\tilde{\zeta}\leq\tilde{P}_{{\mathcal{G}}_{t}}\}}\frac{\tilde{P}_{w}}{\tilde{P}_{{\mathcal{G}}_{t}}}+\mathds{1}_{\{\tilde{\zeta}>\tilde{P}_{{\mathcal{G}}_{t}}\}}\frac{\tilde{P}_{w}}{1-\tilde{P}_{{\mathcal{G}}_{t}}}$
@@
A.5 More Experimental Results
A.5.1 Additional Simulation Studies
Complementary to Figure 6, Figure 7 also shows that the sum test tends to outperform the higher criticism when m is relatively small.

### v2

@@ 5 Language Model Experiments
Setup. We evaluate our watermarking method against three schemes discussed above: (1) the Gumbel-max watermark based on exponential minimal sampling by Aaronson and Kirchner (2023), (2) the soft green/red list watermark by Kirchenbauer et al. (2023a) with a moderate text distortion hyperparameter $\delta=1$, and (3) the Distribution-Preserving watermark (DiPmark) by Wu et al. (2023) with their recommended watermark hyperparameter $\alpha=0.45$. For all methods where applicable, we set $\gamma=0.5$ to control the green list size such that $|{\mathcal{G}}|=\gamma|{\mathcal{W}}|$. For a fair comparison, our scheme uses both the green tokens and the red tokens for detection as described in Section 3.2. Following Fernandez et al. (2023), for all methods, we score tokens based on the uniqueness of the full $(k+1)$-tuple and compute $p$-values using the exact null distribution rather than $z$-scores, and flag a text as watermarked if the $p$-value is lower than $0.01$, which ensures a false positive rate to be less than $0.01$. To ensure fair comparison of text quality, we implement repeated context masking for all methods (Wu et al., 2023).
@@
To evaluate the text distortion induced by the watermark, we use the S-BERT similarity score to compare the semantic similarity between the generated text of the same model with and without watermarking (Reimers and Gurevych, 2019); the higher the similarity score, the less the distortion. We also report the true positive rate (TPR) (i.e., the fraction of correctly identifying a watermarked response among all watermarked samples) before and after attacks, including a purely substitution-based attack (TPR aug.) and a paraphrasing attack (TPR para.). [^10] Following Fernandez et al. (2023), we simulate the substitution attack by randomly and independently replacing each token with probability $0.1$. For the paraphrasing attack, we apply the Context-aware Lexical Substitution (Yang et al., 2022), which introduces a BERT-based infill model for generating contextually appropriate lexical substitutions (Devlin et al., 2019) with paraphrasing ratio$=0.3$. In addition, we investigate the speculative decoding setup to assess the robustness of different watermarks where the goal is to modify watermarked texts so as to improve their quality. As discussed earlier, this setup better approximates the human-language model interaction. Here, we choose Phi-3-medium-4k-instruct (14B parameters) as the target model. [^11] The source code is available at https://github.com/Xieyangxinyu/Debiasing-Watermarks-for-Large-Language-Models-via-Maximal-Coupling. We conduct comprehensive ablation studies with identical experimental configurations, also including Meta’s LLaMA model family (Dubey et al., 2024). Since both model families produce similar results, we focus on the Phi-3 family in the following sections for brevity and present the complete results in the online Appendix.
Table 1: Comparison of the S-BERT similarity scores and the true positive rates (TPR) among four watermarking schemes: (1) the Gumbel-max watermark (Aaronson and Kirchner, 2023), (2) the green/red list watermark (Kirchenbauer et al., 2023a), (3) the DiPmark method Wu et al. (2023), and (4) our proposed one.
@@
TPR para. 0.9950 0.4350 0.7500 0.9250
Comparison against Baselines. Table 1 reports the S-BERT similarity scores and the TPR among the four watermarking schemes. Although the Gumbel-max watermark is theoretically unbiased with the assumption of perfect pseudorandomness (Aaronson and Kirchner, 2023), experimental results show that the deterministic decoder can distort the generated text more severely than the stochastic ones. Indeed, when using the preceding $k$ tokens as the seed for the pseudorandom variables $\xi$ during decoding, perfect pseudorandomness is compromised in practice, as evidenced by the consistently lower S-BERT scores. Meanwhile, when comparing against the green/red list soft watermark and DiPmark, our method often leads to a more competitive TPR when similar S-BERT scores are achieved. This trend holds also under substitution-based and context-aware paraphrasing attacks, as reflected in the TPR aug. and TPR para. metrics.
@@
Text Repetition. We quantitatively evaluate the repetitiveness issue discussed in Section 5 by analyzing the proportion of repeated $k$-token sequences in the generated text. The “Repeated” column in Table 2 reports the average percentage of times that repeated context masking is applied within each generated text, which directly reflects how frequently $k$-grams are repeated during generation. While repeated context masking was introduced by Hu et al. (2023); Dathathri et al. (2024) to mitigate quality degradation, our results show that the Gumbel-max approach still produces notably more repetitive text compared to other methods. For example, with $k=2$ on the FinQA dataset, Gumbel-max exhibits an 18.61% token repetition rate versus 16.77-17.80% for other methods. This suggests that in practice, deterministic dependence on prior $k$-tokens in methods like Gumbel-max can bias the generation toward repetitive patterns that persist even with repeated context masking. In contrast, our stochastic decoder’s additional randomness in sampling helps reduce such repetitive patterns while maintaining watermark effectiveness.
Speculative Decoding. Table 3 presents a comparison of the TPR between the Gumbel-max watermark (Aaronson and Kirchner, 2023) and our method under the speculative decoding setup. This setup, where a draft model proposes tokens that may be rejected by the target model, naturally mirrors how users iteratively refine generated text - making targeted, non-independent edits to improve quality rather than random token substitutions. The average rejection rates indicate the proportion of watermarked tokens that are discarded during this process. While both approaches achieve comparable TPR, the lower rejection rate of our method can translate to improved user experience - users need to make fewer revisions to achieve their desired output compared to the Gumbel-max approach.
Comparison of Test Statistics. In Table 4, we experiment with the two different choices of test statistics: the sum test and HC. Our experiments confirm the insights from Section 4: the sum test consistently achieves higher TPR than HC across all configurations. This aligns with our simulation findings that the sum test outperforms HC when number of tokens used to detect the watermark is relatively small ($\leq 10^{3}$). For longer generated texts, we expect HC to become increasingly competitive with the sum test, as suggested by our simulation results.
Table 3: Comparison of the true positive rates (TPR) and rejection rates between the Gumbel-max watermark and our proposed scheme under the speculative decoding setup.
@@ 6 Conclusions
Acknowledgment
This work was supported in part by the National Key R&D Program of China (2022YFA1007900), NSF DMS-2310679, a Meta Faculty Research Award, Wharton AI for Business, and the National Natural Science Foundation of China (12271013,72342004). This work utilized computational resources provided by the Argonne Leadership Computing Facility and was also based upon work supported by Laboratory Directed Research and Development (LDRD) funding from Argonne National Laboratory, under Contract No. DE-AC02-06CH11357. The authors would like to thank two anonymous referees for their constructive comments that helped improve the presentation of the paper.
Disclosure Statement
The authors report there are no competing interests to declare.
@@ A.1 Unbiased Watermarking Schemes and Related Works
Several alternative methods have been proposed to design provably unbiased watermarking schemes. The main idea involves modifying the decoding strategy of the language model decoder to create watermark signals without altering the token distributions. The first and widely recognized method adapts the exponential minimal sampling, as introduced by Aaronson and Kirchner (2023). This method selects the next token $w_{t}$ directly by computing $w_{t}={\mathcal{S}}(\bm{P}_{t},\zeta_{t}):=\operatorname{argmax}_{w\in{\mathcal{W}}}\log U_{w}/P_{w}$ where $\zeta_{t}=(U_{1},\ldots,U_{|{\mathcal{W}}|})$ is the random variable that consists of $|{\mathcal{W}}|$ i.i.d. copies of $U[0,1]$ and is part of the watermark key. This method is unbiased when true randomness is used for generating $\zeta_{t}$ and the proof involves the Gumbel trick (Gumbel, 1948; Fernandez et al., 2023; Li et al., 2024). However, a common practice to share the watermark key is to seed a pseudorandom number generator for $U$ with the prior $k$ tokens $w_{(t-k)},...,w_{(t-1)}$ (discussed in Section 3.3). Notice that upon fixing the $U$, the decoder is deterministic, unlike the standard decoding methods we use in this paper. When the same $k$-gram appears multiple times in the generated text, this deterministic decoder will generate the same token following these $k$-grams, leading to repetitive texts and degraded text quality (Christ et al., 2024; Kuditipudi et al., 2023). To mitigate this, Wu et al. (2023) propose to apple repeated context masking, which prevents the watermark from being applied on step $t$ if the same $k$-gram has been used to watermark previously. This method has later been adapted by Hu et al. (2023); Dathathri et al. (2024). However, the deterministic nature of this decoder can bias the generation toward repetitive patterns that persist even with repeated context masking, as we demonstrate in Section 5. Several theoretical analyses and variants of this method have been proposed (Fernandez et al., 2023; Zhao et al., 2024), but no existing work completely addresses this issue.
Another method is the inverse transform sampling strategy (Christ et al., 2024; Kuditipudi et al., 2023; Hu et al., 2023; Li et al., 2024). The idea is to first sample a standard uniform random variable $\zeta_{1:n}=(\zeta_{1},...,\zeta_{n})$, and then map $\zeta_{t}$ to the next token. Although the randomness from $\zeta_{t}$ can guarantee the unbiasedness of the token generation, designing an effective test to detect watermarks with this method is challenging: Kuditipudi et al. (2023) showed that their correlation-based detection method yields weaker power than the exponential minimal sampling method, even when perfect randomness is applied, while the detection method proposed by Hu et al. (2023) involves solving an optimization problem and requires hyperparameter tuning to achieve competitive power.
@@ 2. $H_{1}$:
$\Lambda(\zeta_{1:n})=\frac{\prod_{t=1}^{m}\mathbb{1}_{\zeta_{j}\in\left[0,P_{t,{\mathcal{G}}_{t}}\right]}/P_{t,{\mathcal{G}}_{t}}}{\prod_{t=1}^{m}\mathbb{1}_{\zeta_{t}\in\left[0,1\right]}}=(1-\delta)^{-m}\mathbb{1}_{\zeta_{(m)}\leq 1-\delta}$
where $\mathbb{1}$ denotes the indicator function. This likelihood-ratio test provides the decision rule as follows: for some critical value $c$,
@@ 1. $H_{1}^{(\mathrm{mix})}$:
A.3.3 Suboptimality of the Order Statistics.
@@
As observed earlier, we can assume without loss of generality that $P_{t,{\mathcal{G}}_{t}}=P_{{\mathcal{G}}}=m^{-r}$ for all $t$. By Lemma A.1, it suffices to show that the Hellinger affinity between $U[0,1]$ and $(1-\varepsilon_{m})U[0,1]+\varepsilon_{m}U[0,P_{{\mathcal{G}}}]$ behaves asymptotically as $1+o(1/m)$. Let $f(y)$ and $g(y)$ be the densities of $U[0,1]$ and $U[0,P_{{\mathcal{G}}}]$ respectively. As $f(y)=1$ and $g(y)=P_{{\mathcal{G}}}^{-1}\mathbb{1}_{x\in[0,P_{{\mathcal{G}}}]}$ on $[0,1]$, the Hellinger affinity is given by
@@
$=\varepsilon_{m}^{2}\left(\mathbb{E}_{0}p(G)^{-2}\mathbb{1}_{x\in[0,P_{{\mathcal{G}}}]}-1\right)$
@@
$\mathbb{E}_{0}L_{m}\to-\infty\quad\text{and}\quad\frac{\operatorname{Var}_{0}(L_{m})}{[\mathbb{E}_{0}L_{m}]^{2}}\to 0$
@@
$\operatorname{Var}_{1}s=m\operatorname{Var}_{1}\zeta_{1}=O(m)$
@@
$F_{m}(x)=\frac{1}{m}\sum_{i=1}^{m}\mathbb{1}_{\zeta_{i}\leq x},$
@@
$=\sqrt{m}\cdot\left(\frac{\frac{1}{m}\sum_{i=1}^{m}\mathbb{E}_{1}\mathbb{1}_{\zeta_{i}\leq x}-x}{\sqrt{x(1-x)}}\right)$
@@
A.4 Extension to the Soft Green/Red List Watermark
Lemma A.4.
A.5 Speculative Decoding
@@
A.5.1 Connections to Maximal Coupling
@@
$\mathbb{P}[\texttt{Accept}]=\mathbb{P}_{w\sim\bm{Q}}[\zeta\cdot Q_{w}>P_{w}]=\sum_{w\in{\mathcal{W}}}Q_{w}\cdot\min(\frac{P_{w}}{Q_{w}},1)=\sum_{w\in{\mathcal{W}}}\min(P_{w},Q_{w})=p$
@@
A.5.2 Speculative Decoding as Post-Processing
@@
$Q_{w}=\mathbb{1}_{\{w=\operatorname{argmax}_{w^{\prime}\in{\mathcal{W}}}\log U_{w^{\prime}}/\tilde{P}_{w}^{\prime}\}}$
@@
$Q_{w}=\mathbb{1}_{\{\tilde{\zeta}\leq\tilde{P}_{{\mathcal{G}}_{t}}\}}\frac{\tilde{P}_{w}}{\tilde{P}_{{\mathcal{G}}_{t}}}+\mathbb{1}_{\{\tilde{\zeta}>\tilde{P}_{{\mathcal{G}}_{t}}\}}\frac{\tilde{P}_{w}}{1-\tilde{P}_{{\mathcal{G}}_{t}}}$
@@
A.6 More Experimental Results
A.6.1 Additional Simulation Studies
Complementary to Figure 6, Supplementary Figure 7 also shows that the sum test tends to outperform the higher criticism when m is relatively small.